# Byam: Fixing Breaking Dependency Updates with Large Language Models

### v2

@@ 9 Conclusion
Acknowledgements.
This work was supported by the CHAINS project funded by Swedish Foundation for Strategic Research (SSF), the WebInspector project funded by Swedish Research Council (VR), as well as by the Wallenberg Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation. Some computation was enabled by resources provided by the National Academic Infrastructure for Supercomputing in Sweden (NAISS).
10 Declarations
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper. All data and code related to this paper is available at https://github.com/chains-project /bacardi

### v3

@@ 9 Conclusion
Declarations
Funding: This work was supported by the CHAINS project funded by Swedish Foundation for Strategic Research (SSF), the WebInspector project funded by Swedish Research Council (VR), as well as by the Wallenberg Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation. Some computation was enabled by resources provided by the National Academic Infrastructure for Supercomputing in Sweden (NAISS).
Conflict of Interest: None
Ethical approval The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper. All data and code related to this paper is available at https://github.com/chains-project /bacardi
Author Contributions Frank Reyes contributed to the experimental design, data collection, data analysis, and writing. May Mahmoud contributed to the experimental design, data collection, data analysis, and writing. Federico Bono contributed to the experimental design and data collection. Sarah Nadi contributed to data analysis and writing. Benoit Baudry contributed to data analysis and writing. Martin Monperrus contributed to data analysis and writing.
Generative AI: Generative AI was not used for the generation of any part of the content in this paper or fordata analysis. Grammarly, a tool that uses AI, was used for spell checking, grammar correction, and improving writing clarity.
Clinical Trial Number Not applicable